Find the general solution of the Lagrange's equation 2yzp+zxq=3xy?


1
Expert's answer
2022-02-14T17:41:02-0500

The auxiliary equations are

"\dfrac{dx}{2yz} = \dfrac{dy}{zx} = \dfrac{dz}{3xy}"


Taking the first two ratios,

"\begin{aligned}
\dfrac{dx}{2yz} &= \dfrac{dy}{zx}\\
\dfrac{dx}{2y} &= \dfrac{dy}{x}\\
x~dx &= 2y~dy\\
\dfrac{x^2}{2} &= 2\dfrac{y^2}{2}+\dfrac{u}{2}\\
\dfrac{x^2}{2} - y^2 &= \dfrac{u}{2}\\
x^2 - 2y^2 &= u\\
\end{aligned}"


Taking the last two ratios,

"\begin{aligned}
\dfrac{dy}{zx} &= \dfrac{dz}{3xy}\\
\dfrac{dy}{z} &= \dfrac{dz}{3y}\\
3y~dy&= z~dz\\
\dfrac{3y^2}{2} &= \dfrac{z^2}{2} + \dfrac{v}{2}\\
3y^2 -z^2&=v\\
\end{aligned}"


The general solution of the equation is,

"\begin{aligned}
\phi(u,v)&=0\\
\phi(x^2-2y^2,3y^2-z^2)&=0
\end{aligned}"